### v2

# Limited Parallelization in Gate Operations Leads to Higher Space Overhead and Lower Noise Threshold

Abstract
Decoherence of qubits during gate operations in the error correction phase of fault-tolerant computation is significant since, in practice, gate operations are neither instantaneous nor are they sufficiently parallelized. Thus, in practice, the overhead of a fault-tolerant scheme should depend on the level of parallelization, the speed of gate operations, and the rate of decoherence. In this paper, we obtain a lower bound on the required space overhead for any fault-tolerant technique when the decoherence process is Markov and the errors are either erasure or depolarization. An implication of this bound is that the noise or decoherence threshold, i.e., the noise beyond which no reliable memory (circuit) can be realized, decreases with a decrease in the level of parallelization in gate operations. Also, the noise threshold approaches zero as the level of parallelization approaches zero, implying that any fault-tolerant implementations must strive for sufficient parallelization in gate operations.
Fault-tolerant quantum computation happens in two alternating phases: a computation phase, and an error correction phase [1, 2, 3]. In most fault-tolerance models, qubit decoherence and gate errors during the computation phase are modeled together as a noise operator, but the decoherence of qubits during the error correction phase is generally ignored [1, 2, 3, 4, 5]. However, in practice, qubit decoherence in the error correction phase can be significant. This is because in the error correction phase, even if the measurements are instantaneous and accurate, the gate operations (for correction) on qubits after the measurements require non-negligible time [6, 7, 8]. During this time, qubits interact with the bath and decohere. Clearly, if the gate operations during error corrections can be fully parallelized, qubit decoherence would be minimized. However, the level of parallelization in the current technologies is often limited due to issues like cross-talk [9, 10, 11].
In this paper, we want to understand the interplay between the rate or strength of qubit decoherence and the level of parallelization in gate operations, and their impact on fault-tolerance. In particular, we want to characterize the required space overhead in terms of the rate (strength) of qubit decoherence and the amount of parallelization in gate operations during error correction. It is known that for obtaining a lower bound on the space overhead of quantum circuits of a certain depth, one can obtain a lower bound on the space overhead of maintaining a quantum memory for a time proportional to the circuit depth [4]. Here the proportionality constant is the total time taken by a computation phase followed by an error correction phase. Henceforth, we mostly restrict our discussion to fault-tolerant quantum memories, but the results extend to circuits.
Our main contribution is the derivation of a new lower bound on the space overhead requirement for a quantum memory when there is qubit decoherence and the parallelization in gate operations is limited during error correction. The bound on required overhead depends explicitly on the amount of parallelization in gate operations and the rate of decoherence, and is substantially more than the known lower bounds [4, 5]; for example, the bound $\frac{1}{Q(\mathcal{N})}$. Here $Q(\mathcal{N})$ is the quantum capacity of the quantum channel $\mathcal{N}$ which represents noise or errors [12, 13]. The observed increase in the space overhead is the penalty paid for the lack of parallelization in gate operations. Further, for any non-zero decoherence rate, we prove that quantum states stored using any code is unrecoverable after linear (in the number of physical qubits) number of error correction epochs, if the amount of parallelization in gate operations is below a threshold. This threshold depends on the decoherence rate (strength or probability). To the best of our knowledge, this is the first work to analytically characterize the effect of (lack of) gate parallelization on fault-tolerance overhead, which is an important issue for any reasonable size quantum circuit [9, 10, 11].
Model and Main Result
Consider a quantum memory that stores a quantum state of $l$ logical qubits using $n$ physical qubits. This memory is periodically error corrected after a time $\bar{t}$ from the end of the previous error correction phase. An error correction phase is of duration $\bar{t}_{corr}$ and in this time the number of parallel $2$-qubit gate operations is $k=n\frac{\alpha}{2}$ for $10<\alpha<$ 1. Here, $\alpha$ is the maximum level of parallelization of $2$-qubit gate operations that can be achieved with negligible cross-talk [9, 10, 11]. In an error correction epoch, there is no priority among the uncorrected qubits for gate operations.
We allow measurements to be instantaneous and both measurements and gate operations to be error-free. Thus, we consider qubit decoherence during and in between error corrections as the only source of error. Since we are interested in a lower bound on the space overhead, this simplification is fine.
We consider two well-known models for qubit decoherence: erasure and depolarising [12, 13]. The erasures (or depolarizations) are assumed to be i.i.d. across qubits and Markov in time, i.e., probability of erasure of a physical qubit by time $t$ is of the form $1-\exp(-\kappa t)$, for some $\kappa>0$, known as the decoherence rate [12]. Thus, the probability that a qubit, which was fine after the previous error correction phase, gets erased (or depolarizes to $\frac{I}{2}$) in between two error correction phases is $1-\exp(-\kappa\bar{t})$.
Errors due to decoherence happen during the error correction phase as well. Since we are interested in a lower bound on space overhead, we allow the errors to be more conducive to correction. We assume that the decoherence for the error correction phase happens at the beginning of the phase, so that it can be measured and corrected. Thus, we consider a modified error model where errors happen only between two error correction epochs, but the probability of error, $p=1-\exp(-\kappa(\bar{t}+\bar{t}_{corr}))$ depends on the duration of the correction phase, $\bar{t}_{corr}$.
Our main results are the following lower bounds on space overhead and thresholds for gate parallelization.
Theorem 1.
For keeping the above memory with $\alpha<p$ error corrected with a non-vanishing probability for a time duration $n(\bar{t}+\bar{t}_{corr})$ against the above erasure error model, any quantum error correcting code must have $n\geq\frac{l~{}p}{2\alpha-p}$ for $\alpha>\frac{p}{2}$. Moreover, no quantum error correcting code can keep the memory corrected for a time $n(\bar{t}+\bar{t}_{corr})$ if, the level of parallelization, $\alpha<\frac{p}{2}$.
For the above depolarising error model, the lower bound on space overhead is $n\geq\frac{l}{Q_{\mbox{dep}}(\frac{p-\alpha}{p})}$ for $\alpha>\frac{2p}{3}$, where $Q_{\mbox{dep}}(\beta)$ is the quantum capacity of i.i.d. depolarising channel when the probability of depolarization to $\frac{I}{2}$ is $\beta$. Further, no code can keep the memory corrected for a time $n(\bar{t}+\bar{t}_{corr})$ if $\alpha<\frac{2p}{3}$.
This theorem implies that a space overhead more than $\frac{1}{Q(\mathcal{N})}=\frac{1}{1-2p}$ for $p<\frac{1}{2}$, is necessary when errors are erasures and $\alpha<p(1-p)$. This increase in overhead compared to the existing lower bounds [4, 5] captures the cost paid for limitations in speed and parallelization in gate operations. The parallelization threshold also gives a tighter lower bound on noise threshold for erasure errors: fault-tolerant memory (circuits) of arbitrary retention time (depth) are not possible if $p>\min(2\alpha,\frac{1}{2})$. The same threshold on $p$ for depolarizing errors is $\min(\frac{3}{2}\alpha,\frac{1}{3})$.
The above theorem is proved using the following steps. First, we show that with probability almost one, the number of uncorrected qubits after $n\frac{p-\alpha}{\alpha}+1$ phases is more than $n(\frac{p-\alpha}{p}-\epsilon)$ for any $\epsilon>0$, and the errors are uniformly distributed. Then, we derive a lower bound on the space overhead for any quantum code when the number of errors are $n\frac{p-\alpha}{p}$ and the errors are uniformly distributed. This is obtained by relating the overhead required for correcting uniformly distributed $n\gamma$ errors, for $\gamma>0$, to the overhead required for correcting i.i.d. errors with probability $\gamma$. We obtain the threshold on the level of parallelization by deriving the maximum $\gamma$ beyond which no code can correct uniformly distributed $n\gamma$ errors.
Implications and Discussion
As mentioned above, an important implication of the above theorem is a reduction in the noise threshold and an increase in the necessary space overhead when there is a lack of parallelization in gate operations. This leads to a serious requirement on the technological designs: fault-tolerance is impossible if the amount of parallel operations $k$ does not scale linearly with $n$, i.e., $\alpha$ must be $\Omega(1)$. This is because the noise threshold is linear in $\alpha$ for erasure and depolarising errors.
In almost all implementations of gate-based quantum circuits or memory, gate operations take significant time and are not always sufficiently parallelized. The theorem presented here gives a guideline for such implementations on the maximum allowable rate (or minimum required overhead) of the employed quantum error correcting code for a given noise strength and a given amount of parallelization. The theorem leads to a lower bound on the minimum amount of parallelization necessary for a given decoherence rate. Further, it also provides insights into designing $\bar{t}$ for quantum memories for a given $\bar{t}_{corr}$, $\kappa$ and $\alpha$. Here, we discuss a few of those cases.
Naturally, $\bar{t}_{corr}$ cannot be less than the time required for one gate operation, $\bar{t}_{g}$. Let us consider the case $\bar{t}_{corr}=\bar{t}_{g}$ and at most $n\frac{\alpha_{g}}{2}$ number of $2$-qubit parallel gate operations are possible in $\bar{t}_{g}$. Thus, from the theorem, if errors are erasures, we must have $\alpha_{g}>\frac{1-\exp(-\kappa(\bar{t}+\bar{t}_{g}))}{2}$. This implies that a positive $\bar{t}$ and hence, an error corrected quantum memory is not possible, if the parallelization $\alpha_{g}$ is not more than $\frac{1-\exp(-\kappa\bar{t}_{g})}{2}$. This implies that when $\kappa t_{g}\ll 1$, $\alpha_{g}$ must grow by $\frac{1}{2}$ of a unit for every unit increase in $\kappa t_{g}$.
For quantum circuits $\bar{t}=\bar{t}_{g}$ and hence, we need $\alpha_{g}>\frac{1-\exp(-2\kappa\bar{t}_{g})}{2}$, implying that when $\kappa t_{g}\ll 1$, $\alpha_{g}$ must grow equally for every unit for unit increase in $\kappa t_{g}$.
I Proof of the Main Result
After an error correction epoch, each correct qubit undergoes decoherence with probability $p=1-\exp(-\kappa(\bar{t}+\bar{t}_{corr}))$. On the other hand, the qubits that were erased or depolarized after the previous error correction stay that way. During the error correction phase, the measurements are assumed to be accurate and are assumed to give the correct location of the errors. Since our goal is a lower bound on space overhead, this assumption is fine. Note that in the correction phase, at most $k=n\frac{\alpha}{2}$ $2$-qubit gate operations can be performed. Hence, even if the error locations are known, at most $2k=n\alpha$ qubits can be corrected. This leads to the following Markov dynamics for the number of errors.
I-A Markov Chain of Errors
The evolution of the number of errors is modeled as a Markov process as follows. Let $X_{t}$ denote the number of corrupted (decohered) qubits at the start of time epoch $t$. Initially, there are no corrupted qubits ($X_{0}=0$). At the start of time epoch $t$, due to interactions with the environment, each qubit that is not already decohered is acted on by a noise operator (e.g., erasure or depolarizing noise), making it decohere with probability $p$. We assume that the noise acts independently on each qubit, so that the number of qubits that decohere at time epoch $t$ is given by a $\operatorname{Binomial}(n-X_{t},p)$ random variable, which we denote as $Y_{t}$. Then, in the correction phase, we can correct $f_{t}$ of the qubits, where $f_{t}=f(X_{t}+Y_{t})=\min\{X_{t}+Y_{t},n\alpha\}$. The resulting quantum state at the start of time epoch $t+1$ has $X_{t+1}=X_{t}+Y_{t}-f_{t}$ decohered qubits, and the sequence $\{X_{t}\}_{t\in\mathbb{N}}$ forms a Markov chain.
I-B Preliminaries
In this section, we will state some results that will be used throughout the paper. First, we look at the familiar concentration bounds for binomial random variables.
Lemma 1.
(Concentration bounds) Let $X\sim\operatorname{Binomial}(n,p)$. Then,
$\mathds{P}[X\geq n(p+\epsilon)]$ $\leq\exp{(-2n\epsilon^{2})},\text{ and}$
$\mathds{P}[X\leq n(p-\epsilon)]$ $\leq\exp{(-2n\epsilon^{2})}.$
Proof.
By definition, $X$ can be written as a sum of $n$ independent $\operatorname{Bernoulli}(p)$ random variables, each of which is $1/2-$subgaussian (by Hoeffding’s lemma). Therefore, $X$ is $\sqrt{n}/2-$subgaussian, and we get the concentration bounds by applying Chernoff’s inequality and using the subgaussian property [14]. ∎
Next, we mention an important property of the Markov chain $\{X_{t}\}$, related to the monotonicity of $\mathds{P}[X_{t+1}\geq k\mid X_{t}=x]$ with respect to $x$; if $x$ is more, then intuitively this probability should be higher.
Lemma 2.
Let $h_{k}(x)=\mathds{P}[X_{t+1}\geq k\mid X_{t}=x]$. Then, for all $k\in[0,n]$, $h_{k}(\cdot)$ is an increasing function.
Proof.
We have $h_{k}(x)=\mathds{P}[x+Y_{t}-f(x+Y_{t})\geq k\mid X_{t}=x]$. Consider the function $g(z)=z-f(z)$. From the definition of $f$, we observe that $g$ is non-decreasing, and hence for each $k\in[0,n]$ there exists a unique $z_{k}$ (given by $\inf_{z}\{g(z)\geq k\}$) such that $g(z)\geq k\Longleftrightarrow z\geq z_{k}$. Thus, we can write $h_{k}(x)=\mathds{P}[x+Y_{t}\geq z_{k}\mid X_{t}=x]=\mathds{P}[Y_{t}\geq z_{k}-x\mid X_{t}=x]=\mathds{P}[Z\geq z_{k}-x]$, where $Z\sim\operatorname{Binomial}(n-x,p)$.
We can rewrite this probability as $h_{k}(x)=\mathds{P}[n-x-Z\leq n-x-(z_{k}-x)]=\mathds{P}[Z^{\prime}\leq n-z_{k}]$, where $Z^{\prime}=n-x-Z\sim\operatorname{Binomial}(n-x,1-p)$. We may think of $Z^{\prime}$ as a sum of $n-x$ independent $\operatorname{Bernoulli}(1-p)$ random variables. It is now evident that if $x$ increases, then $n-x$ decreases, and thus $h_{k}(x)=\mathds{P}[Z^{\prime}\leq n-z_{k}]$ increases (not necessarily strictly, as $z_{k}-x$ could be negative), because $n-z_{k}$ is a constant independent of $x$. ∎
As a corollary, we have the following extension of the previous lemma.
Lemma 3.
Let $h_{k}^{(m)}(x)=\mathds{P}[X_{t+m}\geq k\mid X_{t}=x]$, where $m\geq 1$. Then, for all $k\in[0,n]$, $h_{k}^{(m)}(\cdot)$ is an increasing function. Further, this implies that $\mathds{P}[X_{t}\geq k]$ is increasing in $t$.
Proof.
We prove by induction on $m$. We have already proved the result for $m=1$ in lemma 2. Now, suppose the result is true for $m=j$. Consider
$\mathds{P}[X_{t+j+1}\geq k\mid X_{t}=x]$ $=\sum_{y=0}^{n}\mathds{P}[X_{t+j+1}\geq k\mid X_{t+j}=y,X_{t}=x]\cdot\mathds{P}[X_{t+j}=y\mid X_{t}=x]$
$=\sum_{y=0}^{n}\mathds{P}[X_{t+j+1}\geq k\mid X_{t+j}=y]\cdot\mathds{P}[X_{t+j}=y\mid X_{t}=x].$
Denote $a_{y}=\mathds{P}[X_{t+j+1}\geq k\mid X_{t+j}=y]$ and $b_{y}=\mathds{P}[X_{t+j}=y\mid X_{t}=x]$. Then, we have
$\mathds{P}[X_{t+j+1}\geq k\mid X_{t}=x]=a_{0}b_{0}+a_{1}b_{1}+\ldots+a_{n}b_{n}$
$=a_{0}(b_{0}+b_{1}+\ldots+b_{n})+(a_{1}-a_{0})(b_{1}+\ldots+b_{n})+\ldots+(a_{n}-a_{n-1})b_{n}$
$=a_{0}\mathds{P}[X_{t+j}\geq 0\mid X_{t}=x]+(a_{1}-a_{0})\mathds{P}[X_{t+j}\geq 1\mid X_{t}=x]\ldots+(a_{n}-a_{n-1})\mathds{P}[X_{t+j}\geq n\mid X_{t}=x].$
By the induction hypothesis, each $\mathds{P}[X_{t+j}\geq y\mid X_{t}=x]$ is increasing in $x$, and we also have $a_{0}\leq a_{1}\ldots\leq a_{n}$ by lemma 2. Thus, the above sum is a non-negative linear combination of increasing functions, and hence $\mathds{P}[X_{t+j+1}\geq k\mid X_{t}=x]$ is increasing in $x$. By induction, the result is true for all $m\geq 1$.
Now, consider for any $t\geq 0$,
$\mathds{P}[X_{t+1}\geq k]$ $=\sum_{x=0}^{n}\mathds{P}[X_{t+1}\geq k\mid X_{1}=x]\cdot\mathds{P}[X_{1}=x]$
[…]
$=\mathds{P}[X_{t+1}\geq k\mid X_{1}=0]$
$=\mathds{P}[X_{t}\geq k\mid X_{0}=0]=\mathds{P}[X_{t}\geq k],$
and hence $\mathds{P}[X_{t}\geq k]$ is increasing in $t$. ∎
I-C Hitting Time Bound
Suppose we have a quantum error correcting code that can correct up to $n\beta$ errors. Then, it is a crucial requirement that $X_{t}\leq n\beta$ after a sufficiently long time, as otherwise we cannot reliably store the system in memory. The quantities that we are most interested in analyzing are $\mathds{P}[X_{t}>n\beta]$, and $\mathbb{E}\big{[}\tau\big{]}$, where $\tau=\inf_{t}\{X_{t}>n\beta\}$ denotes the first time that the number of decohered qubits is more than $n\beta$. In this paper, we only focus on the former.
We will prove that if $\beta$ is smaller than a certain threshold, then $\{X_{t}>n\beta\}$ happens with high probability for all $t>n\beta$. Before proving this result, we state a helpful lemma.
Lemma 4.
Suppose $\beta<\frac{p-\alpha}{p}$, and let $c=\frac{(p-\alpha-p\beta)^{2}}{1-\beta}$. Then, for all $x\leq n\beta$, we have
$\mathds{P}[X_{t+1}>x\mid X_{t}=x]\geq 1-\exp(-2nc).$ (1)
Proof.
Suppose $X_{t}=x$. Recall that, conditioned on $X_{t}=x$, $Y_{t}\sim\operatorname{Binomial}(n-x,p)$ is the number of qubits that decohere at time $t$. Clearly, if $Y_{t}>n\alpha$, we have $X_{t+1}>x$ because at most $n\alpha$ qubits are corrected in the correction phase. Therefore,
$\mathds{P}[X_{t+1}>x\mid X_{t}=x]\geq\mathds{P}[Y_{t}>n\alpha\mid X_{t}=x]=\mathds{P}[\operatorname{Binomial}(n-x,p)>n\alpha].$ (2)
Since $x\leq n\beta$, it follows that $\mathds{P}[\operatorname{Binomial}(n-x,p)>n\alpha]\geq\mathds{P}[\operatorname{Binomial}(n-n\beta,p)>n\alpha]$. Further, we have $(n-n\beta)p=np-n\beta p>n\alpha$, hence from lemma 1 we can lower bound $\mathds{P}[\text{Bin}(n-n\beta,p)>(n-n\beta)(p-\epsilon)=n\alpha]$, where $\epsilon=p-\frac{\alpha}{1-\beta}$. We get
$\mathds{P}[\operatorname{Binomial}(n-x,p)>n\alpha]\geq\mathds{P}[\operatorname{Binomial}(n-n\beta,p)>n\alpha]\geq 1-\exp{(-2nc)},$ (3)
where $c=\frac{(p-\alpha-p\beta)^{2}}{(1-\beta)}$. Thus,
$\mathds{P}[X_{t+1}>x\mid X_{t}=x]=\mathds{P}[X_{t+1}\geq x+1\mid X_{t}=x]\geq 1-\exp{(-2nc)}.$ (4)
We now state the hitting time bound.
Theorem 2.
Suppose $\beta<\frac{p-\alpha}{p}$, and let $c=\frac{(p-\alpha-p\beta)^{2}}{1-\beta}$. Then,
$\mathds{P}[X_{t}>n\beta]\geq\big{(}1-\exp(-2nc)\big{)}^{n\beta+1},$ (5)
for all $t\geq n\beta+1$.
Proof.
We first show that $\mathds{P}[X_{t}\geq t]\geq\big{(}1-\exp(-2nc)\big{)}\cdot\mathds{P}[X_{t-1}\geq t-1]$ whenever $12\leq t\leq n\beta+$ 1. Consider
$\mathds{P}[X_{t}\geq t]$ $\geq\mathds{P}[X_{t}\geq t,X_{t-1}\geq t-1]$ (6)
$=\sum_{x\geq t-1}\mathds{P}[X_{t}\geq t\mid X_{t-1}=x]\cdot\mathds{P}[X_{t-1}=x]$ (7)
[…]
$=\big{(}1-\exp(-2nc)\big{)}\cdot\mathds{P}[X_{t-1}\geq t-1].$ (10)
Since $\mathds{P}[X_{1}\geq 1]=\mathds{P}[X_{1}\geq 1\mid X_{0}=0]\geq 1-\exp(-2nc)$, we have from the above recursion,
$\mathds{P}[X_{t}\geq t]\geq\big{(}1-\exp(-2nc)\big{)}^{t},$ (11)
for all $12\leq t\leq n\beta+$ 1. In particular, substituting $t=n\beta+1$ gives
$\mathds{P}[X_{n\beta+1}\geq n\beta+1]\geq\big{(}1-\exp(-2nc)\big{)}^{n\beta+1},$ (12)
and finally, from lemma 3 it follows that $\mathds{P}[X_{t}\geq n\beta+1]\geq\big{(}1-\exp(-2nc)\big{)}^{n\beta+1}$ for all $t\geq n\beta+1$, as desired. ∎
I-D Proof of Theorem 1
This proof builds on the bound in Theorem 2. Note that for large circuits (large $n$), Theorem 2 implies that with probability almost one, the number of uncorrected errors at time $n\beta+1$ would be at least $n\beta$, for $\beta<\frac{p-\alpha}{p}$. Therefore, we must necessarily have $\beta\geq\frac{p-\alpha}{p}$ in order to ensure that a quantum code that can correct up to $n\beta$ errors can recover the quantum state after a sufficiently long time.
Since the errors are i.i.d. and the gate operations do not prioritize among errors, these $n\beta$ errors are uniformly distributed across $n$ qubits. To prove that memory under erasure cannot be recovered after $n\beta+1$ phases using any quantum code with overhead less than $\frac{p}{2\alpha-p}$, we connect recovery under uniform errors to that under i.i.d. errors.
Consider an i.i.d. quantum erasure channel with probability of erasure $\beta-\delta$. Suppose that there is a code with rate strictly more than $1-2\beta$, that can correct uniform $n\beta$ erasures. For any $\delta>0$, and for all sufficiently large $n$, the number of erasures in i.i.d. erasure channel is upper-bounded by $n\beta$ with probability $1-\exp(-2n\delta^{2})$, by lemma 1. Further, by the property of i.i.d. Bernoulli coin toss, given that a certain number of erasures have happened, the erasures are uniformly distributed. Thus, the code for uniform errors can recover a state with probability $1-\exp(-2n\delta^{2})$ over the i.i.d. $\beta-\delta$ erasure channel. This implies that we have a code with rate strictly more than $1-2\beta$ for an i.i.d. $\beta-\delta$ erasure channel. By taking $\delta$ arbitrarily close to zero, we see that the capacity result of i.i.d. erasure channels is violated. Thus, by contradiction, we get that $n\beta$ uniform erasures cannot be corrected using a code with rate more than $1-2\beta$. Therefore, any code that can correct $n\beta$ uniform erasures must have overhead of at least $\frac{1}{1-2\beta}\geq\frac{p}{2\alpha-p}$.
A similar argument gives the result for depolarising noise. The thresholds for $\alpha$ are obtained by looking into the probability of errors of the respective channels at which the quantum capacities vanish.
II Conclusion
In this paper, we investigated the basic aspects of the limitations imposed by the lack of parallelization in gate operations. The above results and discussions show that this limitation can lead to severe issues like high space overhead and vanishing noise thresholds. These results lead to some necessary guidelines for designing practical systems. The interplay between gate parallelization and issues like gate failures, correlation in errors, and non-Markovian noise, which are of practical consequence, remain to be explored.
Acknowledgement
The authors would like to thank SERB, India for support through CRG/2023/005345.